OpenStudy (atomicreaper):

Find the roots of x^2/4=2x-10 .

OpenStudy (jango_in_dtown):

x^2/4=2x-10 implies x^2=8x-40 implies x^2-8x+40=0 roots \[4\pm2\sqrt{6}i\]